A Bibliography of Publications of Jim Grundy

Jim Grundy

Department of Computer Science The Australian National University

Canberra ACT 0200 AUSTRALIA Tel: ?n/a?

FAX: +61 (0)6 249-0010

E-mail: Jim.Grundy@cs.anu.edu.au (Internet)

25 November 2011 Version 1.15

Abstract

This bibliography records publications of Jim Grundy.

Title word cross-reference

11th [GN98b].

9th [vWGH96a].

August [vWGH96b].

Browsable [Gru96a].

Calculational [BGv96a]. Calculator [BGL^{+}97]. Conference [vWGH96a].

emerging [GN98b]. EVES [Gru91c]. Exchange [Gru96d].

Format [Gru96a].

Hierarchical [Gru96c]. Higher [vWGH96a]. HOL [Gru91a].

Inference [Gru91d].

Library [Gru91a]. Logic [Gru93c]. Logics [vWGH96a].

m [Gru91c]. m-EVES [Gru91c]. Method [Gru93a].

October [GN98b]. Order [vWGH96a].

pair [Gru91a]. Predicative [Gru93b]. Presentation [Gru96a]. Proceedings [vWGH96a]. Program [Gru93a]. Programming [Gru93b]. Proof [Gru96a]. Proofs [GL96]. Proving [vWGH96a].

Reasoning [Gru96c]. Record [GL96]. Recording [GL97]. Refinement [Gru92]. Report [Gru91c].

September [GN98b]. Storage [Gru96d]. Structured [BGv96a]. Supplementary [vWGH96a]. Support [BGL^{+}97]. Survey [Gru93b]. System [Gru91d].

Theorem [vWGH96a]. Theorems [Gru96d]. Three [Gru93c]. Three-Valued [Gru93c]. Tool [Gru92]. TPHOLs’96 [vWGH96a]. TPHOLs’98 [GN98b]. Transformational [Gru96c]. trends [GN98b]. Trustworthy [Gru96d].

Valued [Gru93c].

window [Gru91b].

References

Butler:1997:RCP

[BGL^{+}97] Michael Butler, Jim Grundy, Thomas Långbacka, Rimvydas Rukšėnas, and Joakim von Wright. The refinement calculator: Proof support for program refinement. In Lindsay Groves and Steve Reeves, editors, Formal Methods Pacific’97: Proceedings of FMP’97, Discrete Mathematics & Theoretical Computer Science, pages 40– 61. Springer-Verlag, Wellington, New Zealand, July 1997.

Grundy:1996:SCA

[BGv96a] Ralph Back, Jim Grundy, and Joakim von Wright. Structured calculational proof. Joint Computer Science Technical Report TR-CS-96-09, Department of Computer Science, The Australian National University, Canberra ACT 0200, Australia, November 1996.

Grundy:1996:SCT

[BGv96b] Ralph Back, Jim Grundy, and Joakim von Wright. Structured calculational proof. TUCS Technical Report 65, Turku Centre for Computer Science, Lemminkäisenkatu 14, FIN-20520 Turku, Finland, November 1996.

Grundy:1996:TBR

[GL96] Jim Grundy and Thomas Långbacka. Towards a browsable record of HOL proofs. TUCS Technical Report 7, Turku Centre for Computer Science, Lemminkäisenkatu 14, FIN-20520 Turku, Finland, May 1996.

Grundy:1997:RHP

[GL97] Jim Grundy and Thomas Långbacka. Recording HOL proofs in a structured browsable format. In Michael Johnson, editor, Algebraic Methodology and Software Technology: 6th International Conference, AMAST’97, volume 1349 of Lecture Notes in Computer Science, pages 567–571. Springer-Verlag, Sydney, Australia, December 1997.

Grundy:1998:TPH

[GN98a] Jim Grundy and Malcolm Newey, editors. Theorem proving in higher order logics: 11th international conference, TPHOLs’98, Canberra, Australia, September 27 – October 1, 1997: proceedings, volume 1479 of Lecture Notes in Computer Science.

Springer-Verlag, Berlin, 1998. ISBN 3-540-64987-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96T655 1998.

Grundy:1998:TPE

[GN98b] Jim Grundy and Malcolm Newey, editors. Theorem proving in higher order logics: emerging trends: 11th international conference, TPHOLs’98, Canberra, Australia, September 27 – October 1, 1997: supplementary proceedings, number 98-08 in Joint Computer Science Technical Reports. Department of Computer Science, The Australian National University, Canberra ACT 0200, Australia, 1998. ISBN 0-7315-4800-0.

Grundy:1991:HPL

[Gru91a] Jim Grundy. The HOL pair library. In The HOL System, volume Libraries. SRI International, Cambridge Research Center, Millers Yard, Mill Lane, Cambridge CB2 1RQ, England, 2.01 edition, July 1991.

Grundy:1991:HWL

[Gru91b] Jim Grundy. The HOL window library. In The HOL System, volume Libraries. SRI International, Cambridge Research Center, Millers Yard, Mill Lane, Cambridge CB2 1RQ, England, 2.01 edition, July 1991.

Grundy:1991:RME

[Gru91c] Jim Grundy. Report on m-EVES. ERL Research Report ERL-0545-RR, Information Technology Division, Defence Science & Technology Organisation, PO Box 1500, Salisbury SA 5108, Australia, March 1991.

Grundy:1991:WIH

[Gru91d] Jim Grundy. Window inference in the HOL system. In Myla Archer, Jeffrey J. Joyce, Karl N. Levitt, and Phillip J. Windley, editors, Proceedings of the International Tutorial and Workshop on the HOL Theorem Proving System and its Applications, pages 177–189. ACM-SIGDA, IEEE Computer Society Press, University of California at Davis, August 1991.

Grundy:1992:WIT

[Gru92] Jim Grundy. A window inference tool for refinement. In Cliff B. Jones, B. Tim Denvir, and Roger C. F. Shaw, editors, Proceedings of the 5th Refinement Workshop, Workshops in Computing, pages 230–254. BCS-FACS, Springer-Verlag, Lloyd’s Register, London, January 1992.

Grundy:1993:MPR

[Gru93a] Jim Grundy. A Method of Program Refinement. Phd thesis, technical report 318, Computer Laboratory, University of Cambridge, New Museums Site, Pembroke Street, Cambridge CB2 3QG, England, November 1993.

Grundy:1993:PPS

[Gru93b] Jim Grundy. Predicative programming — A survey. In Dines Bjørner, Manfred Broy, and Igor V. Pottosin, editors, Formal Methods in Programming and Their Applications: Proceedings of the International Conference, volume 735 of Lecture Notes in Computer Science, pages 8– 25. Springer-Verlag, Novosibirsk, Russia, June 1993.

Grundy:1993:TVL

[Gru93c] Jim Grundy. A three-valued logic for refinement. In Dines Bjørner, Manfred Broy, and Igor V. Pottosin, editors, Formal Methods in Programming and Their Applications: Proceedings of the International Conference, volume 735 of Lecture Notes in Computer Science, pages 26–42. Springer-Verlag, Novosibirsk, Russia, June 1993.

Grundy:1996:ABF

[Gru96a] Jim Grundy. A browsable format for proof presentation. Mathesis Universalis, 1(2), Spring 1996.

Grundy:1996:BFP

[Gru96b] Jim Grundy. A browsable format for proof presentation. In Christoffer Gefwert, Pekka Orponen, and Jouko Seppänen, editors, Logic, Mathematics and the Computer — Foundations: History, Philosophy and Applications, volume 14 of Symposium Series, pages 171–178. Finnish Artificial Intelligence Society, Helsinki, Finland, June 1996.

Grundy:1996:THR

[Gru96c] Jim Grundy. Transformational hierarchical reasoning. The Computer Journal, 39(4):291– 302, May 1996. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Grundy:1996:TSE

[Gru96d] Jim Grundy. Trustworthy storage and exchange of theorems. TUCS Technical Report 1, Turku Centre for Computer Science, Lemminkäisenkatu 14, FIN-20520 Turku, Finland, April 1996.

Wright:1996:SPI

[vWGH96a] Joakim von Wright, Jim Grundy, and John Harrison, editors. Supplementary Proceedings of the 9th International Conference on Theorem Proving in Higher Order Logics: TPHOLs’96, number 1 in TUCS General Publications. Turku Centre for Computer Science, Turku, Finland, August 1996. ISBN 951-650-781-6. ISSN 1239-1905.

Wright:1996:TPH

[vWGH96b] Joakim von Wright, Jim Grundy, and John Harrison, editors. Theorem proving in higher order logics: 9th International Conference, TPHOLs’96, Turku, Finland, August 26–30, 1996: proceedings, volume 1125 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1996. CODEN LNCSD9. ISBN 3-540-61587-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96I577 1996.